OpenStudy (anonymous):

Beni's uncle is building a deck for her house that is proportional to the twenty- feet by thirty- feet deck he built for his house. The scale factor between his deck and her deck is zero point five. Which of the following could be a measurement for Beni's deck? PLEASE HALP

OpenStudy (anonymous):

well since the scale factor is 0.5 or 1/2 we can set up a proportion for each measurement:\[\frac{ 1 }{ 2 } = \frac{ x }{ 20 } and \frac{ 1 }{ 2 } = \frac{ x }{ 30 }\]. If you cros multiply you get 2x = 20 and 2x = 30. Now solver for x by dividing each term by 2 and you get x = 10ft and x = 15ft

OpenStudy (anonymous):

@ditzzzz omg thank you so much i was so lost.